OpenStudy (anonymous):

A 75-kg sprinter accelerates from rest to a speed of 11.0 m/s in 5.0 s.? a) Calculate the mechanical work done by the sprinter during this time. I know W=0.5mv^2 = 4538 J b)Calculate the average power the sprinter must generate P=W/t= 908W c) If the sprinter converts food energy to mechanical energy with an efficiency of 25%, at what average rate is he burning Calories? (I dont get this question) Thank You!

OpenStudy (anonymous):

calory is a form of energy. rate of burning calory is power. efficiency=25% (given) power emitted or generated= 908W therefore, power generated/(rate of burning of calory)=25/100 so, rate of burning of calory= 4*908 = 3632W